If a car starts it journey at 11 am with velocity = 60km/h then an0ther car after one hour starts its journey with velocity = 80km/h. When does 2nd car cross 1st?
a. 1 pm
b. 1:50 pm
c. 2 pm
d. 2:30pm

Answer by Theo(13342)

Rate * Time = Distance
That's the formula you use.
They will both have traveled the same distance when the first car crosses the second car.
The Equation for the first car is:
60 * T = D
The Equation for the second car is:
80 * (T-1) = D
T is the number of hours the first car travels.
T-1 is the number of hours the second car travels because the second car starts 1 hour after the first car and so is traveling 1 hour less when they cross each other.
Your 2 equations are:
60 * T = D
80 * (T-1) = D
since they are both equal to D, we can set them equal to each other to get:
60 * T = 80 * (T-1)
Simplify to get:
60 * T = 80 * T - 80
subtract 60 * T from both sides of the equation and add 80 to both sides of the equation to get:
80 = 20 * T
divide both sides of the equation by 20 to get:
4 = T
T is equal to 4 hours.
T-1 is equal to 3 hours.
The first car travels 4 hours at 60 kilometers per hour to travel a total distance of 240 kilometers.
The second car travels 3 hours at 80 kilometers per hour to travel a total distance of 240 kilometers
They will cross each other 240 miles from the start6ing point.
This will happen 4 hours after the first car starts.
The first car starts at 11:00 am.
The crossing occurs as 3:00 pm.
Unfortunately, that's not one of your answer selections.
I double-checked the calculations and they look good.
?????